OpenStudy (anonymous):

Matrix help \left[\begin{matrix}4 & 5 \\ 0 & -1\end{matrix}\right] +2 \left[\begin{matrix}1 & 8 \\ -4 & -6\end{matrix}\right]

OpenStudy (anonymous):

\[\left[\begin{matrix}4 & 5 \\ 0 & -1\end{matrix}\right] +2 \left[\begin{matrix}1 & 8 \\ -4 & -6\end{matrix}\right]\]